At retirement a man was informed by SSNIT that his total contribution with interest over his working years was 300,000.if the man is anticipated to die in 20 years time how much will he receive monthly over the 20 years before he dies if a discount rate of 12% per annum compounded monthly is used by SSNIT
1
Expert's answer
2020-12-20T17:58:04-0500

"p=\frac{300000}{20 \cdot 12\cdot (1+\frac{12}{12\cdot 100})^{20}}=1024.43."